Chris, a small business owner, has identified a store site in a shopping mall in Atlanta. The mall owner offered him to alternatives: one location with 10,000 square feet and the other with 15,000 square feet. If he chooses the larger space, he estimates that the store will make a $200,000 annual profit when the economy is strong. However, if the economy is normal or poor, the profit will be only $100,000 or $40,000 respectively. With the smaller store, he estimates the annual profit at $150,000 with a good economy, $100,000 with a normal one and $50,000 with a poor one. Economists predict that there is a 10% probability of a boom, an 80% chance of a normal economy, and a 10% chance of a poor one in the coming year. Which alternative is better for Chris to maximize his store profit?

a.) Choose the smaller space.

b.) Choose the large space.
Answer #1

Option (b).

Expected profit, Smaller space ($) = 10% x 150,000 + 80% x 100,000 + 10% x 50,000 = 15,000 + 80,000 + 5,000 = 100,000

Expected profit, Larger space ($) = 10% x 200,000 + 80% x 100,000 + 10% x 40,000 = 20,000 + 80,000 + 4,000 = 104,000

Larger space has higher expected profit.